OpenStudy (anonymous):

Write the equation of a line that passes through the point (1,-3) and is parallel to the line 7x + y = -5.

OpenStudy (anonymous):

A. y = -7x + 4 B. y = -4x + 7 C. y = 4x + -7 D. y = 28x

OpenStudy (anonymous):

General form of all linear equations is y=mx + c where m is the gradient and c in the y intercept Rearrange the given equation in this form and you get y = -7x -5 the gradient is -7 for you. For two lines that are parallel, the gradient is the same. If you want to find the equation of a line and you know its gradient and the the point it passes through, you can use the relation y-y1=m(x-x1) where (x1, y1) is the point the line passes through. Just plug in the values you have y-(-3) = -7(x-1) y+3 = -7x+7 y = -7x+4 Answer